The elastic limit of copper is 𝟏. πŸ“ Γ— πŸπŸŽπŸ– π‘΅π’Žβˆ’πŸ . If a copper wire is stretched by a load of mass πŸ–. πŸ“ π’Œπ’ˆ then find the minimum diameter of the wire within the elastic limit.Β 


1
Expert's answer
2021-11-29T11:43:41-0500
"p=4\frac{mg}{\pi d^2}\\1.5\cdot10^8=4\frac{8.5\cdot9.8}{\pi d^2}\\d=8.4\cdot10^{-4}\ m"